OpenStudy (anonymous):

the length of the room is 1 more than 3 times the width. the area of the room is 80 square meters. what is the width of the room?

OpenStudy (johnweldon1993):

Well we know that Area = Length x Width... A = L x W And we know the length is 1 (+1) more than 3 times (3w) the width... so L = 3W + 1 So we plug everything in we know... A = L x W Area = 80 L = 3W + 1 so 80 = (3W + 1)W Distribute that W 80 = 3w^2 + w Subtract the 80 from both sides... \[\large 3w^2 + w - 80 = 0\] Looks like a quadratic to me...can you take this from here?